THE MICROWAVE SPECTRA OF THE LINEAR OC HCCCN, OC DCCCN, AND THE T-SHAPED HCCCN CO_{2} COMPLEXES

\underline{LU KANG}, Department of Natural Sciences, Union College, Barbourville, KY 40906; STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The rotational spectra of OC HCCCN, OC DCCCN, and HCCCN CO_{2} have been recorded using pulsed-jet Fourier transform microwave spectroscopy. Accurate molecular constants have been obtained and agree with the results from IR spectroscopy.^{a}^{,} ^{b} The OC HCCCN is hydrogen-bonded in a linear configuration, and the HCCCN CO_{2} is T-shaped with the CO_{2} molecule at the nitrogen end of the cyanoacetylene. The ^{13}C, the ^{15}N, and the ^{18}O isotopologues for both the OC HCCCN and OC DCCCN complexes have been observed in natural abundance. The fully substituted molecular structure of the OC HCCCN complex has been determined. The results will be compared with those of the corresponding complexes of HCN.

aX. Yang, R. Z. Pearson, and G. Scoles, Chem. Phys. Lett. 204, 145-151 (1993)

bX. Yang, R. Z. Pearson, and G. Scoles, J. Mol. Spectrosc. 180, 1-6 (1996)